# Exploring decay and -delayed neutron emission in exotic 46,47Cl isotopes

### v1

@@ I Introduction
Experimental investigations away from the valley of stability have highlighted unique and interesting properties of exotic nuclei, including neutron halos, neutron skins, and the disappearance of the standard magic numbers. For extremely neutron-rich nuclei, -delayed neutron emission becomes an important pathway in the decay process. With increasing neutron number, the -value for decay increases and, simultaneously, the neutron-separation energy () in the daughter nucleus decreases. This leads to an increase in the population of neutron-unbound states in the allowed decay process. Additionally, for nuclei with a large excess of neutrons, the parent and daughter nuclei often have different ground-state parities due to the valence protons and neutrons occupying opposite-parity shells. As a result, 1-particle, 1-hole () states, which typically lie at high energies, are preferentially populated in allowed decay. This, in part, drives the shifting of the Gamow-Teller strength above the neutron-emission threshold. These states typically decay by emitting a neutron, leading to a daughter nucleus with , though, in principle, decay by -ray emission cannot be completely ruled out. For extremely neutron-rich nuclei, some Gamow-Teller strength can be located above the two-neutron separation energy, which will result in the emission of more than one neutron following decay.
@@ II Experimental Setup
As shown in Fig. 1, particle identification was performed event-by-event using energy loss in two Si PIN detectors and their timing information relative to the plastic scintillator at the A1900 intermediate dispersive image. Downstream of the Si PIN detectors, the secondary beam particles, including the ions of interest, 46,47Cl, were implanted in the 40 strips x 40 strips pixelated Double-Sided Silicon Strip Detector (DSSD) belonging to the Beta Counting System (BCS) [32]. The implantation rate was about 150/s which allowed efficient correlation of implants and decays. An Al degrader was placed upstream to ensure that implants stopped near the middle of the 986-m thick DSSD. A Single-Sided Silicon Strip Detector (SSSD) was installed downstream to veto light particles punching through the DSSD. The BCS was surrounded by 16 HPGe Clover detectors to record -delayed rays. Energy and efficiency calibration was performed using standard -ray sources up to 3.5 MeV. The efficiency was about 5% at 1 MeV. The setup did not include neutron detection, and, hence, information about delayed neutron emission was obtained from delayed -ray transitions instead. The data were collected using the NSCL digital data acquisition system which provides timestamps for each channel, allowing coincidences and correlations to be built in the offline analysis [33].
@@ III.1 Half-life measurements
30 | 109.5 | 129.3 |
The decay curves for 46,47Cl are shown in Figs. 2(a) and 2(b) respectively. These curves were generated by histogramming the time difference between the selected implanted ion (46Cl or 47Cl) and its decay product. The particle was required to be detected within the same pixel as the implant or in one of the neighboring eight pixels of the implant within a time correlation window of 2 s. Each decay curve was fitted with an expression to account for the exponential decay of the parent nucleus and the growth and decay of the daughter nuclei following the Bateman equations. The resulting half-life for 46Cl is 201(16) ms which agrees with the previous measurement of 223(37) ms [35, 36] within uncertainties and is slightly lower than the value of 232(2) ms from Ref. [37] but within 2. The half-life measured for 47Cl in this work is 109.5(25) ms, which is slightly higher than the literature value of 101(6) ms [37], but within their mutual uncertainties. In both cases, the exponential growth and decay of the daughter nuclei using the literature values for their half-lives [34] and an exponential background from long-lived activities were accounted for. Additionally, the 47Cl half-life was determined by gating on the 1553-keV -ray transition in the daughter, 46Ar (Fig. 2(c)). An exponential (parent decay) and a constant (background) were fitted to the curve, yielding a half-life of 109(5) ms, in excellent agreement with the ungated fit.
@@
-delayed -ray spectra correlated with 46,47Cl implants are shown in Figs. 3(a) and (b). Time correlation windows of 250 ms and 100 ms were used for 46Cl and 47Cl, respectively. These time intervals correspond to roughly one half-life for the two cases and highlight decays from the parent nucleus. For 46Cl decay (Fig. 3(a)), the spectrum is dominated by the 1553(2)-keV and 542(1)-keV transitions which are the ground-state decays of the first excited states in the daughter (46Ar) and the daughter (45Ar), respectively. For 47Cl decay (Fig. 3(b)), the spectrum is again dominated by the 1553(2)-keV transition in 46Ar ( daughter), signaling that the -delayed neutron emission channel is stronger than the decay to neutron-bound states. The 1231-keV transition represents the lowest one in 47Ar, the daughter.
@@
- coincidences observed in 47Ar following 47Cl decay are shown in Fig. 4(a). A gate on the known 1231(1)-keV transition () in 47Ar highlights its coincidence with the 1602(2)-keV and 1935(2)-keV transitions. These transitions have not been reported previously, and correspond to new levels at 2833(2) keV and 3166(2) keV, respectively. The ground state of 47Cl is conjectured to be = from the present half-life measurement (see Table 1), and, thus, these states are candidates for positive-parity states populated by allowed Gamow-Teller (GT) transitions. Fig. 4(b) shows known - coincidences in the granddaughter nucleus 47K [38] which follows the decay of 47Ar confirming its production in the relatively weak branch of the decay of 47Cl.
@@
The transition in 46Ar at 1553 keV is the most intense transition observed in both the 46Cl and 47Cl decays (Figs 3(a),(b)). The coincidences observed with the 1553-keV transition in the two cases are shown in Figs. 5(a) and (b). From 46Cl decay, a 1143(3)-keV transition was observed in coincidence with the 1553(2)-keV transition, as seen in Fig. 5(a). This could correspond to the 1140(20)-keV peak from Ref. [22]. 46Cl decay was also reported in a conference proceedings where 2006-keV and 3350-keV rays were observed in coincidence with each other and with the 1553-keV transition [39]. We were able to confirm these coincidence relationships (Fig. 5(b)) and further suggest that the 2009(3)-keV transition feeds into the 3347(3)-keV transition, although the reverse cannot be ruled out, as both transitions are equally intense within the estimated uncertainties. The statistics for 46Cl decay were rather low in this experiment and no further -ray transitions could be identified. The absolute feeding to the observed levels in 46Ar from 46Cl decay and the corresponding log values are given in Table 2.
(keV) | (I) | log |
1553(2) | 49(24)111should be considered an upper limit due to possible unobserved transitions | 8.34(33) [FF] |
2696(2) | 2.6(6) | 9.38() [FF] |
3249(3) | 3.7(8) | 9.10() [FF] |
4900(3) | 0 | |
6909(4) | 7.5(21) | 5.51() [GT] |
@@
From 47Cl decay, several transitions were observed in coincidence with the 1553-keV peak in 46Ar, as seen in Fig. 5(c). Due to low statistics, these transitions were difficult to observe in an ungated spectrum. The 2133(2), 2311(2), 2517(2), and 2699(2)-keV transitions in the present work may correspond to the 2141(3), 2318(3), 2518(2), 2707(2)-keV transitions cited in Ref. [24], although they are somewhat lower in energy. The 2307(13) and 2692(16)-keV rays reported in the inverse kinematics proton scattering of 46Ar in Ref. [23] are similar in energy to the 2311 and 2699-keV transitions seen in the present work. The 3037(2), 3197(3), 3427(3), and 3627(3)-keV transitions are reported for the first time in this work. In Ref. [23] a 3430(26)-keV ray was observed in coincidence with the transition and could correspond to the 3427-keV transition seen here. However, other decays observed in Ref. [23] from that state were not observed in this work, leading us to suggest that they could be distinct. The relative intensities of the rays observed in coincidence with the 1553-keV transition are provided in Table 3 for both 46Cl and 47Cl decay.
47Cl decay | 46Cl decay | ||
@@
1193(1) | 24(9) | 1193(1) | 22(13) |
1230(1) | 91(18) | 1230(1) | 100 |
1525(1) | 63(16) | 1525(1) | 27(16) |
2755(2) | 100 | 2755(2) | 63(31) |
799(1) | 68(20) |
The value for 47Cl is 16040(100) keV, while is 4.0(3) MeV [34]. This allowed us to observe -ray transitions in the daughter, 45Ar, which were also independently populated via the 46Cl channel ( = 7.96(10) MeV). For both cases, clean coincidences with the 542(1)-keV ray (de-excitation of the first excited state to the ground state in 45Ar) were observed. The 46Cl and 47Cl delayed -ray spectra gated on the 542(1)-keV transition are compared in Figs. 6(a) and (b), and the corresponding partial level schemes are also displayed. For 45Ar from 47Cl decay, only decays from positive-parity states above the 542-keV level were observed [30, 39]. For 45Ar populated from 46Cl decay, the rays from the same set of positive-parity states are seen along with a strong 799(1)-keV transition. Based on the 45Cl decay scheme reported in Refs. [30, 39], this ray is assigned to the de-excitation of the state at 1341(1) keV. Relative intensities of the -ray transitions observed in the 542-keV gate are provided in Table 4 for both cases.
@@ IV Discussion
The level scheme of 46Ar, which is the daughter of 46Cl and the daughter of 47Cl was established using - coincidences (Fig. 5) and energy and intensity balance. The partial level schemes with states in 46Ar populated in 46Cl decay (left) and 47Cl decay (right) are shown in Fig. 7. Interestingly, though not a complete surprise, the levels beyond the first observed from 46Cl decay and 47Cl decay are completely different. The low spin states predicted from shell model calculations using the FSU interaction [25] and the interaction [26] are displayed alongside for comparison.
@@
A state was reported at 1745 keV in a prior in-beam -ray spectroscopy experiment following one-proton knockout [40]. The same experiment proposed a state at 2186 keV and states at 1231 keV, 2763 keV, and 3437 keV [40]. These states are indicated by green lines in Fig. 8 and as can be seen they align closely in energy to levels with the same predicted by and FSU interactions. The previously-observed level at 1745 keV nicely corroborates with the shell-model state at 1371 keV from the interaction and 1890 keV state as predicted using the FSU interaction. The interaction provides candidate for the 2188 keV and 2763 keV states in the and states. This would leave open the possibility of the present 2833-keV and 3166-keV levels tentatively being the (2281 KeV) and (2550 keV) states though somewhat lower in energy. The equivalent states calculated using the FSU interaction are at 2365 keV () and 3074 keV ().
@@
The ground-state configuration of 47Cl () is with a filled neutron shell and a hole in the proton orbital. The states with proton excitations into the shell are energetically too high in 46Ar to be reached by GT decay (i.e conversion of one of the neutrons into a proton). Instead, a GT transformation changes one of the neutrons in either the or orbitals to a proton, filling the orbital, as shown by the arrow in the top cartoon. This leaves a hole in the or orbital resulting in a positive parity state of 47Ar. The ground state configuration of 47Ar () is with a full neutron shell and an unpaired neutron in the orbital. We can assume that the (log = 5.3) and (log = 5.5) states in 47Ar predicted at 4.95 MeV and 5.18 MeV, respectively, are populated in this allowed GT decay from the ground state of 47Cl. With equal to 3.667(3) MeV, these states lie above the neutron separation threshold in 47Ar.

### v2

@@ I Introduction
Experimental investigations away from the valley of stability have highlighted unique and interesting properties of exotic nuclei, including neutron halos, neutron skins, and the disappearance of the standard magic numbers. For extremely neutron-rich nuclei, -delayed neutron emission becomes an important pathway in the decay process. With increasing neutron number, the value for decay increases and, simultaneously, the neutron-separation energy () in the daughter nucleus decreases. This leads to an increase in the population of neutron-unbound states in the allowed decay process. Additionally, for nuclei with a large excess of neutrons, the parent and daughter nuclei often have different ground-state parities due to the valence protons and neutrons occupying opposite-parity shells. As a result, 1-particle, 1-hole () states, which typically lie at high energies, are preferentially populated in allowed decay. This, in part, drives the shifting of the Gamow-Teller strength above the neutron-emission threshold. These states typically decay by emitting a neutron, leading to a daughter nucleus with , though, in principle, decay by -ray emission cannot be completely ruled out. For extremely neutron-rich nuclei, some Gamow-Teller strength can be located above the two-neutron separation energy, which will result in the emission of more than one neutron following decay.
@@ II Experimental Setup
As shown in Fig. 1, particle identification was performed event-by-event using energy loss in two Si PIN detectors and their timing information relative to the plastic scintillator at the A1900 intermediate dispersive image. Downstream of the Si PIN detectors, the secondary beam particles, including the ions of interest, 46,47Cl, were implanted in the 40 strips x 40 strips pixelated Double-Sided Silicon Strip Detector (DSSD) (active area of about 40 mm x 40 mm) belonging to the Beta Counting System (BCS) [32]. An Al degrader was placed upstream before the DSSD to ensure that implants were stopped near the middle of the 986-m thick DSSD. Furthermore, the straggling through the degrader ensured that a large number of the DSSD pixels were illuminated. The implantation rate was about 150/s spread over the greater than 1000 pixels of the DSSD. This rate was sufficient to allow the decay of the implanted radioactive ion (half-lives typically 100s of ms or less) before a second implantation in the same or neighboring pixel, enabling clean correlations between the chosen implant and its decay products. As the DSSD was 1 mm thick, the detection efficiency was only 50-60%. A Single-Sided Silicon Strip Detector (SSSD) was installed downstream to veto light particles punching through the DSSD.
The BCS was surrounded by 16 HPGe Clover detectors to record -delayed rays. Energy and efficiency calibration was performed using standard -ray sources up to 3.5 MeV. The efficiency was about 5% at 1 MeV. The setup did not include neutron detection, and, hence, information about delayed neutron emission was obtained from delayed -ray transitions instead. The data were collected using the NSCL digital data acquisition system which provides timestamps for each channel, allowing coincidences and correlations to be built in the offline analysis [33].
@@ III.1 Half-life measurements
| 30 | 109.5 | 129.3 | |||
The decay curves for 46,47Cl are shown in Figs. 2(a) and 2(b) respectively. These curves were generated by histogramming the time difference between the selected implanted ion (46Cl or 47Cl) and its correlated decay event. The particle was required to be detected within the same pixel as the implant or in one of the neighboring eight pixels of the implant within a time correlation window of 2 s. Each decay curve was fitted with an expression which included the exponential decay of the parent nucleus and the growth and decay of the daughter nuclei (0n) following the Bateman equations. In both cases, literature values of half-lives of daughter nuclei were used [34]. As the n daughters namely, 46Ar () and 45Ar (), are very long lived they were not explicitly included. Instead, an additional slowly falling exponential functional with free parameters was added to account for long-lived activities and constant background for each isotope. The resulting half-life for 46Cl is 201(16) ms which agrees with the previous measurement of 223(37) ms [35, 36] within uncertainties and is slightly lower than the value of 232(2) ms from Ref. [37] but within 2. The half-life measured for 47Cl in this work is 109.5(25) ms, which is slightly higher than the literature value of 101(6) ms [37], but within their mutual uncertainties. Additionally, the 47Cl half-life was determined by gating on the 1553-keV -ray transition in the daughter, 46Ar (Fig. 2(c)). An exponential (parent decay) and a constant (background) were fitted to the curve, yielding a half-life of 109(5) ms, in excellent agreement with the ungated fit.
@@
-delayed -ray spectra correlated with 46,47Cl implants are shown in Figs. 3(a) and (b). Time correlation windows of 250 ms and 100 ms were used for 46Cl and 47Cl, respectively. These time intervals correspond to roughly one half-life for the two cases and highlight decays from the parent nucleus. For 46Cl decay (Fig. 3(a)), the spectrum is dominated by the 1553(2)-keV and 542(1)-keV transitions which are the ground-state decays of the first excited states in the daughter (46Ar) and the daughter (45Ar), respectively. For 47Cl decay (Fig. 3(b)), the spectrum is again dominated by the 1553(2)-keV transition in 46Ar ( daughter), signaling that the -delayed neutron emission channel is stronger than the decay to neutron-bound states. The 1231-keV transition is from the decay of the first excited state in 47Ar, the daughter.
@@
- coincidences observed in 47Ar following 47Cl decay are shown in Fig. 4(a). A gate on the known 1231(1)-keV transition () in 47Ar highlights tentative coincidences with the 1602(2)-keV and 1935(2)-keV transitions. These transitions have not been reported previously, and correspond to new levels at 2833(2) keV and 3166(2) keV, respectively. The ground state of 47Cl is conjectured to be = from the present half-life measurement (see Table 1), and, thus, these states are candidates for positive-parity states populated by allowed Gamow-Teller (GT) transitions. Fig. 4(b) shows known - coincidences in the granddaughter nucleus 47K [38] which follows the decay of 47Ar confirming its production in the relatively weak branch of the decay of 47Cl.
@@
The transition in 46Ar at 1553 keV is the most intense transition observed in both the 46Cl and 47Cl decays (Figs. 3(a) and (b)). The coincidences observed with the 1553-keV transition in the two cases are shown in Figs. 5 and 6. From 46Cl decay, a 1143(3)-keV transition was observed in coincidence with the 1553(2)-keV transition, as seen in Fig. 5(a). This could correspond to the 1140(20)-keV peak from Ref. [22]. 46Cl decay was also reported in a conference proceedings where 2006-keV and 3350-keV rays were observed in coincidence with each other and with the 1553-keV transition [39]. We were able to confirm these coincidence relationships (Fig. 5(b)) and further suggest that the 2009(3)-keV transition feeds into the 3347(3)-keV transition, although the reverse cannot be ruled out, as both transitions are equally intense within the estimated uncertainties. The absolute intensities of the 1553, 1141, 1695, 2009, and 3347-keV were estimated to be 70(13)%, 2.6(1)%, 3.6(1.5)%, 7(2)%, and 7.5(2.5)% respectively. The statistics for 46Cl decay were rather low in this experiment and no further -ray transitions could be identified. The absolute feeding to the observed levels in 46Ar from 46Cl decay and the corresponding log values are given in Table 2.
| (I) | log | |||
| keV | % | GT | 1FU | 1FNU |
| 1553(2) | 56(28)111should be considered an upper limit due to possible unobserved -ray transitions | - | - | 5.59() |
| 2696(2) | 2.6(7) | - | 9.35() | 6.75() |
| 3248(3) | 3.7(10) | - | 9.10() | 6.51() |
| 4900(3) | 0 | - | - | - |
| 6909(4) | 7.5(25) | 5.51() | - | - |
@@
From 47Cl decay, several transitions were observed in coincidence with the 1553-keV peak in 46Ar, as seen in Fig. 6. Due to low statistics, some of these transitions were difficult to observe in an ungated spectrum. The 2133(2), 2311(2), 2517(2), and 2699(2)-keV transitions in the present work may correspond to the 2141(3), 2318(3), 2518(2), 2707(2)-keV transitions cited in Ref. [24], although they are somewhat lower in energy. The 2307(13) and 2692(16)-keV rays reported in the inverse kinematics proton scattering of 46Ar in Ref. [23] are similar in energy to the 2311 and 2699-keV transitions seen here. The 3037(2), 3197(3), 3427(3), and 3627(3)-keV transitions are reported for the first time in this work. For these transitions with low intensities, effort was made to check them in various time correlation windows and gates (when possible) to avoid possible misidentification, though some ambiguity cannot be ruled out. In Ref. [23] a 3430(26)-keV ray was observed in coincidence with the transition and could correspond to the 3427-keV transition seen here. However, other decays observed in Ref. [23] from that state were not observed in this work, leading us to suggest that they could be distinct. The relative intensities of the rays observed in 46Ar which were possible to observe in a singles spectrum are provided in Table 3 for both 46Cl and 47Cl decay.
| 46Cl decay | 46Cl decay | ||
@@
| 542(1) | 100(12) | 542(1) | 100(12) |
| 799(1) | 30(4) | 1193(1) | 14(3) |
| 1193(1) | 13(2) | 1230(1) | 222intensity could not be extracted due to mixing with the 1231 keV transition from 47Ar. |
| 1230(1) | 35(5) | 1525(1) | 18(3) |
| 1341(1) | 10(3) | 2755(2) | 40(6) |
| 1525(1) | 20(4) | - | - |
| 2755(2) | 6.5(15) | - | - |
The value for 47Cl is 15790(200) keV, while is 4.0(3) MeV [40]. This allowed us to observe -ray transitions in the daughter, 45Ar, which were also independently populated via the 46Cl channel ( = 7.96(10) MeV [40]). For both cases, clean coincidences with the 542(1)-keV ray (de-excitation of the first excited state to the ground state in 45Ar) were observed. The 46Cl and 47Cl delayed -ray spectra gated on the 542(1)-keV transition are compared in Figs. 7(a) and (b), and the corresponding partial level schemes are also displayed. For 45Ar from 47Cl decay, only decays from positive-parity states above the 542-keV level were observed [30, 39]. For 45Ar populated from 46Cl decay, the rays from the same set of positive-parity states are seen along with a strong 799(1)-keV transition. Based on the 45Cl decay scheme reported in Refs. [30, 39], this ray is assigned to the de-excitation of the state at 1341(1) keV. The direct decay from the 1341-keV state is also observed as seen in Fig. 3(a). The relative intensities of the -ray transitions observed in 45Ar are provided in Table 4 for both 46Cl and 47Cl decay.
@@ IV Discussion
The level scheme of 46Ar, which is the daughter of 46Cl and the daughter of 47Cl was established using - coincidences (Fig. 5 and Fig. 6) and energy and intensity balance. The partial level schemes with states in 46Ar populated in 46Cl decay (left) and 47Cl decay (right) are shown in Fig. 8. Interestingly, though not a complete surprise, the levels beyond the first observed from 46Cl decay and 47Cl decay are completely different. The low spin states predicted from shell model calculations using the FSU interaction [25] and the interaction [26] are displayed alongside for comparison.
@@
A state was reported at 1745 keV in a prior in-beam -ray spectroscopy experiment following one-proton knockout [42]. The same experiment proposed a state at 2188 keV and states at 1231 keV, 2761 keV, and 3438 keV [42]. These states are indicated by green lines in Fig. 9 and as can be seen they align closely in energy to levels with the same predicted by and FSU interactions. The previously-observed level at 1745 keV nicely corroborates with the shell-model state at 1371 keV from the interaction and 1890 keV state as predicted using the FSU interaction. The interaction provides candidate for the 2188 keV and 2763 keV states in the and states. This would leave open the possibility of the present 2833-keV and 3166-keV levels tentatively being the (2281 KeV) and (2550 keV) states though somewhat lower in energy. The equivalent states calculated using the FSU interaction are at 2365 keV () and 3074 keV ().
@@
The ground-state configuration of 47Cl () is with a filled neutron shell and a hole in the proton orbital. The states with proton excitations into the shell are energetically too high in 46Ar to be reached by GT decay (i.e conversion of one of the neutrons into a proton). Instead, a GT transformation changes one of the neutrons in either the or orbitals to a proton, filling the orbital, as shown by the arrow in the top cartoon. This leaves a hole in the or orbital resulting in a positive parity state of 47Ar. The ground state configuration of 47Ar () is with a full neutron shell and an unpaired neutron in the orbital. We can assume that the (log = 5.3) and (log = 5.5) states in 47Ar predicted at 4.95 MeV and 5.18 MeV, respectively, are populated in this allowed GT decay from the ground state of 47Cl. With equal to 3.667(3) MeV [40], these states lie above the neutron separation threshold in 47Ar.